OpenStudy (dukelaxfan):

The temperature of a chemical reaction ranges between −10 degrees Celsius and 50 degrees Celsius. The temperature is at its lowest point when t = 0, and the reaction completes 1 cycle during a 6-hour period. What is a cosine function that models this reaction?

OpenStudy (dukelaxfan):

f(t) = −30 cos pi over 3t + 20 f(t) = −60 cos pi over 3t + 30 f(t) = 30 cos 6t + 20 f(t) = 60 cos 6t + 30

OpenStudy (dukelaxfan):

@Directrix @phi @mathmale

OpenStudy (dukelaxfan):

Its "A" right?

OpenStudy (jdoe0001):

hmmm what did you get for the amplitude?

OpenStudy (dukelaxfan):

-30

OpenStudy (jdoe0001):

well, is 30, but negative in this case anyhow, to make it short starts off at -10, goes up to 50, its midline is above the x-axis by 20units, so +20 anyway, yes, it'd be \(y=-30cos\left( \cfrac{\pi }{3}t \right)+20\)

OpenStudy (dukelaxfan):

thanks

OpenStudy (jdoe0001):

yw